OpenStudy (anonymous):

what number must you add to complete the square? x^2+8x=11

OpenStudy (jdoe0001):

let us forget about the 11 for a sec what do you need to complete the square at \(\bf x^2+8x\) ?

hero (hero):

x^2 + 8x + (b/2)^2 = 11 + (b/2)^2

hero (hero):

figure out what b equals...that should help you out

OpenStudy (anonymous):

Thank you for the help ! i really appreciate it !

hero (hero):

Did you figure out the number you're supposed to add to both sides?

OpenStudy (anonymous):

Yes i did ((:

hero (hero):

You mind letting us know so that we know you got it?

OpenStudy (anonymous):

can you help me on this one ?? x^2+4x=3 find the number you must add to complete the square

hero (hero):

I'll help you with that one once you confirm to me that you understood the previous one. Let me know the number you add to both sides. It is possible that you may not have the correct number.

hero (hero):

I'm not trying to doubt your abilities or anything. I just want to make sure that you have it correct.

OpenStudy (anonymous):

i got 4 o: for the other question

hero (hero):

Four is not correct.

hero (hero):

Try to figure out why....And this is exactly what I mean...You assumed 4 was correct, but actually, there's something you have to do after you get 4.

OpenStudy (jdoe0001):

hmmm, did you mean \(\bf 4 \ or \ 4^2 \ \ \ ?\)

OpenStudy (anonymous):

No im saying 4 for the second question iasked, i have a math class online and after i select the answer it tells me whether im correct or not .__. and it said correct

hero (hero):

Which is exactly why you should not post two different questions in the same thread.

OpenStudy (anonymous):

x^2+4x=3 that qanswer i got for that was 4

OpenStudy (anonymous):

Thats why i said second question -______- ?